the question i need help with asks me to solve this equation. 1/x-1/x+4=1/3. i tried and the solution i came up with is 11/3 is this correct. thanks
Answer by checkley77(12844)

IF you mean:
1/x-1/(x+4)=1/3
(x+4-x)/x(x+4)=1/3
4/(x^2+4x)=1/3
x^2+4x=12
x^2+4x-12=0
(x+6)(x-2)=0
x+6=0
x=-6 ans.
x-2=0
x=2 ans.
---------------------
OR:
1/(x-1)/(x+4)=1/3
1/(x-1)*1/(x+4)=1/3
1/(x-1)(x+4)=1/3
1/(x^2+3x-4)=1/3
x^2+3x-4=3
x^2+3x-4-3=0
x^2+3x-7=0
--------------------------
OR:
(1/x)-(1/x)+4=1/3
0+4=1/3
4=1/3 no solution.